A particle mass 10 gm is executing shm with an amplitude of 0.5 m and periodic time of pi/s secThe max value of the force acting on the particle

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

x=Xsinwt         where w=2 pi/T          and T=pi/5

 =.5sin 10t

w=√k/m

so k=w2m

       =102 *10

        =1000   gm/sec2

maximum force=kx

                       =1000*.5  m*gm/sec2

                        =.5 kg m/sec2


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed

solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin


Askiitians Expert Bharath-IITD
14 years ago

Dear Pallavi,

We know for shm motion of the particle  its displacement is given by y= y0 sin ωt  Where  ω = 2π/T =  2π/π = 2 and amplitude

y0 = 0.5 m

thus aceeleration on the particle is given by a=  d2y/d2t =   - y0 * ω2 * sin ωt

Maximum acceleration is amax = - y0 * ω2  = - 2 m/s2

Thus maximum force in magnitude = | m* amax| = 0.02 N

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best  !!!

 


Regards,

Askiitians Experts

Adapa Bharath